# Reverse Thinking Makes LLMs Stronger Reasoners

### v1

@@
Sayna Ebrahimi3, Long Le2, Vincent Perot3, Swaroop Mishra3,
@@
Correspondence to: cychen@cs.unc.edu
@@
Justin Chih-Yao Chen1thanks: Correspondence to: cychen@cs.unc.edu,  Zifeng Wang2, Hamid Palangi2, Rujun Han2, Sayna Ebrahimi3, Long Le2, Vincent Perot3, Swaroop Mishra3, Mohit Bansal1,  Chen-Yu Lee2,  Tomas Pfister2 1UNC Chapel Hill, 2Google Cloud AI Research, 3Google DeepMind
Reverse thinking plays a crucial role in human reasoning. Humans can reason not only from a problem to a solution but also in reverse, i.e., start from the solution and reason towards the problem. This often enhances overall reasoning performance as it enables consistency checks between their forward and backward thinking. To enable Large Language Models (LLMs) to perform reverse thinking, we introduce Reverse-Enhanced Thinking (RevThink), a framework composed of data augmentation and learning objectives. In RevThink, we augment the dataset by collecting structured forward-backward reasoning from a teacher model, consisting of: (1) the original question, (2) forward reasoning, (3) backward question, and (4) backward reasoning. We then employ three objectives to train a smaller student model in a multi-task learning fashion: (a) generate forward reasoning from a question, (b) generate a backward question from a question, and (c) generate backward reasoning from the backward question. Experiments across 12 datasets covering commonsense, math, and logical reasoning show an average 13.53% improvement over the student model’s zero-shot performance and a 6.84% improvement over the strongest knowledge distillation baselines. Moreover, our method demonstrates sample efficiency – using only 10% of the correct forward reasoning from the training data, it outperforms a standard fine-tuning method trained on 10×10$\times$10 × more forward reasoning. RevThink also exhibits strong generalization to out-of-distribution held-out datasets.
@@ 2 Related Work
A large body of research has shown that LLM reasoning can be improved via advanced test-time approaches, such as prompting and aggregation. Representative methods include Chain-of-Thought (CoT) Kojima et al. (2022); Wei et al. (2022) and Self-Consistency Wang et al. (2022), Tree-of-Thought prompting Yao et al. (2024), Self-Reflection Shinn et al. (2024); Madaan et al. (2024); Yao et al. (2022), Multi-agent collaboration Du et al. (2023); Liang et al. (2023); Wang et al. (2023); Lu et al. (2024); Feng et al. (2024); Chen et al. (2023). Several works have been proposed to leverage backward reasoning to verify the chain-of-thought and improve math reasoning Weng et al. (2022); Jiang et al. (2024), while effective, these methods operate on test time, showing moderate improvements compared to other test-time methods like self-consistency Wang et al. (2022). Also, these methods have mostly been developed for mathematical tasks, limiting their generalizability. In contrast, RevThink trains the student model using carefully curated data, enabling it to develop backward reasoning skills in a structured manner. This approach maintains the same test-time efficiency as zero-shot prompting, while delivering greater improvements and generalizing to a broader range of tasks.
@@ 5.1 Main Results
We present our main result in Table 1. First, RevThink demonstrates superior average performance, outperforming all baselines across datasets and models. Compared to the zero-shot performance of the student model, RevThink achieves an average improvement of 12.68% with Mistral and 14.37% with Gemma. Additionally, when compared to SKD and Distill Step-by-Step, which rely on supervised fine-tuning using the correct reasoning chains from the teacher model, RevThink shows substantial improvements of 6.44% to 7.15%. Compared to augmentation-based baselines, RevThink exhibits greater performance gains, particularly in the areas of commonsense reasoning, tabular reasoning and date understanding. While some of these augmentation methods, e.g., Answer Augmentation (AnsAug) are effective for math reasoning, they tend to show less improvements in other domains, suggesting that math, being a more structured domain, scales better with additional data Li et al. (2024); Yuan et al. (2023). In contrast, RevThink consistently improves performance across a variety of reasoning tasks. Later in Table 3, we also show that when evaluating on held-out datasets, RevThink shows larger gains on an out-of-domain math dataset.
@@ 5.2 Additional Analysis
Having demonstrated that RevThink outperforms all baselines with the full training set, we now explore the performance of RevThink and the SKD baseline with varying portions of the training data, denoted by $p\in\{0.1,0.25,0.5,1.0\}$. For instance, when $p=0.1$, we sample 10% of the correct forward reasoning for SKD fine-tuning and apply our data augmentation approach, as described in Section 3, for our fine-tuning. Results shown in Fig. 3 demonstrate that RevThink exhibits strong sample efficiency. Across multiple reasoning tasks, RevThink consistently outperforms SKD at all levels of $p$, even surpassing SKD at $p=1.0$ with only 10% of the data on StrategyQA.
@@
Backward question generation boosts performance, but the full use of our dataset yields the best performance. Recall that each instance in our teacher data is a tuple $(Q^{(i)},R^{(i)}_{f},Q^{(i)}_{b},R^{(i)}_{b})$, consists of the original question, forward reasoning, backward question and backward reasoning. We analyze which combination of components enhances the student model the most. In Fig. 4, we find that learning from all components leads to the best performance. Besides this, we find: (1) only learning from answering backward questions hurts the performance. In Fig. 4, using only $Q_{b}\rightarrow R_{b}$ results in the worst performance and can even fall below the original student model’s zero-shot performance. This is likely because reverse questions are not fully in-domain, and focusing solely on them can lead to distributional shifts, which harms overall performance. (2) learning to generate backward questions can improve the student model. Here we refer to $(Q\rightarrow R_{f})\&(Q\rightarrow Q_{b})$, where in this setting, the student model has two learning objectives: one is to answer the question by generating forward reasoning, and two is to generate the reverse question given the original question. We observe that adding backward question generation to the learning objectives already improves the student model. However, we show that adding $Q_{b}\rightarrow R_{b}$ which enables the student to also learn to answer the reverse question, is the most effective.
@@
RevThink scales positively with model size. We apply RevThink on StrategyQA using models of varying sizes (all models are instruction-tuned). As shown in Fig. 6, our method demonstrates a clear upward trend in accuracy as model size increases. For each model size, applying our method leads to consistent improvements. Notably, Mistral-7B with our method surpasses Mistral-8x22B by 8.36%, despite the latter having 25×25$\times$25 × more parameters. These results underscore that RevThink scales effectively with the model size.
@@
RevThink complements existing methods. With Answer Augmentation (AnsAug) being the best-performing augmentation baseline, we show the complementary strength of RevThink. In Table 4, we compare the performance of AnsAug alone, RevThink alone and their combination. For AnsAug + RevThink, we follow the same augmented dataset and procedure as in Section 3 – with the only difference being sampling another correct reasoning chain for each question from the teacher model – as AnsAug does. We find that while both AnsAug and RevThink yield individual improvements over SKD, their combination results in even greater enhancements, indicating that RevThink effectively complements existing data augmentation methods like AnsAug.
@@ Appendix E Dataset Statistics
SQA Geva et al. (2021) Commomsense MIT 2,061 1,544 229
CSQA Talmor et al. (2019) Commomsense MIT 9,741 6,478 1,140
ARC Clark et al. (2018) Commomsense CC BY-SA 4.0 1,199 1,035 1,172

### v2

@@
Sayna Ebrahimi3, Long T. Le2, Vincent Perot3, Swaroop Mishra3,
@@
Work done as a student researcher at Google Cloud AI Research. Correspondence to: cychen@cs.unc.edu.
@@
Justin Chih-Yao Chen1thanks: Work done as a student researcher at Google Cloud AI Research. Correspondence to: cychen@cs.unc.edu.,  Zifeng Wang2, Hamid Palangi2, Rujun Han2, Sayna Ebrahimi3, Long T. Le2, Vincent Perot3, Swaroop Mishra3, Mohit Bansal1,  Chen-Yu Lee2,  Tomas Pfister2 1UNC Chapel Hill, 2Google Cloud AI Research, 3Google DeepMind
Reverse thinking plays a crucial role in human reasoning. Humans can reason not only from a problem to a solution but also in reverse, i.e., start from the solution and reason towards the problem. This often enhances overall reasoning performance as it enables consistency checks between their forward and backward thinking. To enable Large Language Models (LLMs) to perform reverse thinking, we introduce Reverse-Enhanced Thinking (RevThink), a framework composed of data augmentation and learning objectives. In RevThink, we augment the dataset by collecting structured forward-backward reasoning from a teacher model, consisting of: (1) the original question, (2) forward reasoning, (3) backward question, and (4) backward reasoning. We then employ three objectives to train a smaller student model in a multi-task learning fashion: (a) generate forward reasoning from a question, (b) generate a backward question from a question, and (c) generate backward reasoning from the backward question. Experiments across 12 datasets covering commonsense, math, and logical reasoning show an average 13.53% improvement over the student model’s zero-shot performance and a 6.84% improvement over the strongest knowledge distillation baselines. Moreover, our method demonstrates sample efficiency – using only 10% of the correct forward reasoning from the training data, it outperforms a standard fine-tuning method trained on 10×10$\times$10 × more forward reasoning. RevThink also exhibits strong generalization to out-of-distribution held-out datasets. [^1] Code and data are publicly available at this url.
@@ 2 Related Work
A large body of research has shown that LLM reasoning can be improved via advanced test-time approaches, such as prompting and aggregation. Representative methods include Chain-of-Thought (CoT) Kojima et al. (2022); Wei et al. (2022) and Self-Consistency Wang et al. (2022), Tree-of-Thought prompting Yao et al. (2024), Self-Reflection Shinn et al. (2024); Madaan et al. (2024); Yao et al. (2022), Multi-agent collaboration Du et al. (2023); Liang et al. (2023); Wang et al. (2023); Lu et al. (2024); Feng et al. (2024); Chen et al. (2023). Several works have been proposed to leverage backward reasoning to verify the chain-of-thought and improve math reasoning Weng et al. (2022); Xue et al. (2023); Jiang et al. (2024). While effective, these methods operate on test time, showing moderate improvements compared to other test-time methods like self-consistency Wang et al. (2022). Also, these methods have mostly been developed for mathematical tasks, limiting their generalizability. In contrast, RevThink trains the student model using carefully curated data, enabling it to develop backward reasoning skills in a structured manner. This approach maintains the same test-time efficiency as zero-shot prompting, while delivering greater improvements and better generalizability.
@@ 5.1 Main Results
We present our main result in Table 1. First, RevThink demonstrates superior average performance, outperforming all baselines across datasets and models. Compared to the zero-shot performance of the student model, RevThink achieves an average improvement of 12.68% with Mistral and 14.37% with Gemma. When compared to SKD and Distill Step-by-Step, which rely on supervised fine-tuning using the correct reasoning chains from the teacher model, RevThink shows substantial improvements of 6.44% to 7.15%. RevThink also exhibits greater performance gains compared to augmentation-based baselines, particularly on commonsense reasoning, tabular reasoning and date understanding. While some of these augmentation methods, e.g., Answer Augmentation (AnsAug) are effective for math reasoning, they tend to show less improvements in other domains, suggesting that math, being a more structured domain, scales better with additional data Li et al. (2024); Yuan et al. (2023). In contrast, RevThink consistently improves performance across a variety of reasoning tasks. Later in Table 3, we also show that when evaluating on held-out datasets, RevThink shows larger gains on an out-of-domain math dataset, pointing to its stronger generalizability compared to AnsAug.
@@ 5.2 Additional Analysis
Having demonstrated that RevThink outperforms all baselines with the full training set, we now explore the performance of RevThink and the SKD baseline with varying portions of the training data, denoted by $p\in\{0.1,0.25,0.5,1.0\}$. For instance, when $p=0.1$, we sample 10% of the correct forward reasoning for SKD fine-tuning and apply our data augmentation approach, as described in Section 3, for our fine-tuning. Results shown in Fig. 3 demonstrate that RevThink exhibits strong sample efficiency. Across multiple reasoning tasks, RevThink consistently outperforms SKD at all levels of $p$, even surpassing SKD at $p=1.0$ with only 10% of the data on StrategyQA. Furthermore, while SKD stagnates with varying $p$ on Date Understanding, RevThink shows a clear upward trend in performance when $p$ increases.
@@
Backward question generation alone already boosts performance, but the full use of our dataset yields the best performance. Recall that each instance in our teacher data is a tuple $(Q^{(i)},R^{(i)}_{f},Q^{(i)}_{b},R^{(i)}_{b})$, consists of the original question, forward reasoning, backward question and backward reasoning. We analyze which combination of components enhances the student model the most. In Fig. 4, we find that learning from all components leads to the best performance. Besides this, we find: (1) only learning from answering backward questions hurts the performance. In Fig. 4, using only $Q_{b}\rightarrow R_{b}$ results in the worst performance and can even fall below the original student model’s zero-shot performance. This is likely because reverse questions are not fully in-domain, and focusing solely on them can lead to distributional shifts, which harms overall performance. (2) learning to generate backward questions can improve the student model. Here we refer to $(Q\rightarrow R_{f})\&(Q\rightarrow Q_{b})$, where in this setting, the student model has two learning objectives: one is to answer the question by generating forward reasoning, and two is to generate the reverse question given the original question. We observe that adding backward question generation to the learning objectives already improves the student model. However, we show that adding $Q_{b}\rightarrow R_{b}$ which enables the student to also learn to answer the reverse question, is the most effective.
@@
RevThink scales positively with model size. We apply RevThink on StrategyQA using models of varying sizes (all models are instruction-tuned). As shown in Fig. 6, our method improves each model regarding of its size, i.e., for each model size, applying our method leads to consistent improvements. Even for the largest model, Mistral-8x22B, applying our method still yields a 14.9% improvement. Notably, Mistral-7B with our method surpasses Mistral-8x22B’s zero-shot performance by 8.36%, despite the latter having 25×25$\times$25 × more parameters. These results underscore that RevThink scales effectively with the model size.
@@
RevThink complements existing methods. With Answer Augmentation (AnsAug) being the best-performing augmentation baseline, we show the complementary strength of RevThink. In Table 4, we compare the performance of AnsAug alone, RevThink alone and their combination. For AnsAug + RevThink, we follow the same procedure outlined in Section 3 – with the only difference being sampling another correct reasoning chain for each question from the teacher model – as AnsAug does. We find that while both AnsAug and RevThink yield individual improvements over SKD, their combination results in even greater enhancements, indicating that RevThink effectively complements existing methods like AnsAug.
@@ Appendix E Dataset Statistics
SQA Geva et al. (2021) Commonsense MIT 2,061 1,544 229
CSQA Talmor et al. (2019) Commonsense MIT 9,741 6,478 1,140
ARC Clark et al. (2018) Commonsense CC BY-SA 4.0 1,199 1,035 1,172